# Classical and Quantum KMS States on Spin Lattice Systems

### v1

@@ 1 Introduction
In this framework $j\in\mathbb{Z}_{+}/2$ is interpreted as a semiclassical parameter —in fact, $h:=1/(2j+1)$ is the proper semiclassical parameter— and $j\to\infty$ corresponds to the semiclassical limit. Given a sequence $(\omega_{j})_{j\in\mathbb{Z}_{+}/2}$ of states such that $\omega_{j}\in S(\mathfrak{A}_{j})$, the semiclassical limit is obtained considering the weak*-limit points of the sequence $(\omega_{j}\circ Q_{j})_{j\in\mathbb{Z}_{+}/2}$ of functionals over $\mathfrak{A}_{\infty}$ —each such weak*-limit point defines a state on $\mathfrak{A}_{\infty}$.
@@ 2.2 Berezin SDQ for a single site system
identify a SDQ over a suitably defined bundle of $C^{*}$-algebras $\boxed{B_{\ast}}\subseteq\prod_{j\in\overline{\mathbb{Z}_{+}/2}}B_{j}$, cf. [37, Thm. 8.1].
@@
where $P_{\ell,m}$ denotes the Legendre polynomial of order $\ell,m$. Notice that this choice of normalization is such that $\|Y_{\ell,m}\|_{L^{2}(\mathbb{S}^{2},\mu_{\ell})}=1$, moreover, $\|Y_{\ell,m}\|_{B_{\infty}}\leq 1$, cf. [50, Cor. 2.9]. The set $\{Y_{\ell,m}\}_{\ell,m}$ is a complete orthogonal system for $L^{2}(\mathbb{S}^{2},\mu_{0})$ made by orthogonal but not $L^{2}$-normalized vectors. With this convention we also have [42, §3.6.2]
@@ 2.3 Berezin SDQ on $\Gamma$
$\big{\|}Q_{j}^{\Gamma}\big{(}\{a_{\Lambda},\tilde{a}_{\Lambda}\}_{B_{\infty}^{\Gamma}}\big{)}-i(2j+1)\big{[}Q_{j}^{\Gamma}(a_{\Lambda}),Q_{j}^{\Gamma}(\tilde{a}_{\Lambda})\big{]}\big{\|}_{B_{j}^{\Gamma}}\\ =\big{\|}Q_{j}^{\Lambda}\big{(}\{a_{\Lambda},\tilde{a}_{\Lambda}\}_{B_{\infty}^{\Lambda}}\big{)}-i(2j+1)[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{\Lambda}\tilde{a}_{\Lambda}]\big{\|}_{B_{j}^{\Lambda}}\underset{j\to\infty}{\longrightarrow}0\,,$
@@ 3.2 Semi-classical limit of quantum KMS condition
The proof requires two main ingredients. The first one is a useful characterization of the $(\beta,\delta_{\infty}^{\Gamma})$-KMS classical condition, cf. Lemma 3.2: The latter is inspired by an analogous characterization of $(\beta,\delta_{j}^{\Gamma})$-KMS quantum states, cf. Remark 3.3. The second piece of information is a control of the semiclassical limit of the derivation $\delta_{j}^{\Gamma}$. Specifically, for any finite region $\Lambda\Subset\Gamma$, the DGR condition (3) implies that $(2j+1)\delta_{j}^{\Lambda}\circ Q_{j}^{\Lambda}-Q_{j}^{\Lambda}\circ\delta_{j}^{\Lambda}\underset{j\to\infty}{\longrightarrow}0$ strongly on $\dot{B}_{\infty}^{\Lambda}$. This is a key property which notably holds also for the lattice system on the entire $\Gamma$, cf. Lemma 3.4.
@@
$\lim_{j\to\infty}\Big{\|}(2j+1)\delta_{j}^{\Lambda}(Q_{j}^{\Lambda}(a_{\Lambda}))-Q_{j}^{\Lambda}(\delta_{\infty}^{\Lambda}(a_{\Lambda}))\Big{\|}_{B_{j}^{\Lambda}}=0\qquad\forall a_{\Lambda}\in\dot{B}_{\infty}^{\Lambda}\,.$ (42)
@@
$\lim_{j\to\infty}\Big{\|}(2j+1)\delta_{j}^{\Gamma}(Q_{j}^{\Gamma}(a_{\Lambda}))-Q_{j}^{\Gamma}(\delta_{\infty}^{\Gamma}(a_{\Lambda}))\Big{\|}_{B_{j}^{\Gamma}}=0\,.$ (43)
@@
$\Big{\|}(2j+1)\delta_{j}^{\Gamma}\big{(}Q_{j}^{\Gamma}(a_{\Lambda})\big{)}-Q_{j}^{\Gamma}\big{(}\delta_{\infty}^{\Gamma}(a_{\Lambda})\big{)}\Big{\|}_{B_{j}^{\Gamma}}\\ \leq\sum_{X\cap\Lambda\neq\varnothing}\Big{\|}\frac{2j+1}{i}[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{X}(\varphi_{X})]-Q_{j}^{\Lambda\cup X}\big{(}\{a_{\Lambda},\varphi_{X}\}_{B_{\infty}^{\Lambda}}\big{)}\Big{\|}_{B_{j}^{\Lambda\cup X}}\,.$
We will now prove that the series
$\sum_{X\cap\Lambda\neq\varnothing}\Big{\|}\frac{2j+1}{i}[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{X}(\varphi_{X})]-Q_{j}^{\Lambda\cup X}\big{(}\{a_{\Lambda},\varphi_{X}\}_{B_{\infty}^{\Lambda}}\big{)}\Big{\|}_{B_{j}^{\Lambda\cup X}}\,,$ (44)
converges, moreover, it vanishes in the limit $j\to\infty$. To this avail, let $X\Subset\Gamma$ be such that $X\cap\Lambda\neq\varnothing$: For simplicity we will assume that $\Lambda\subset X$, the general case being treated analogously. We decompose
$\varphi_{X}=\sum_{k=1}^{N_{X}}c_{k,X}\varphi^{k}_{\Lambda}\otimes\varphi^{k}_{X\setminus\Lambda}\,,$
where $N_{X}\in\mathbb{N}$, $c_{j,X}\in\mathbb{R}$, $\varphi^{k}_{\Lambda}\in\dot{B}_{\infty}^{\Lambda}$ and $\varphi^{k}_{X\setminus\Lambda}\in\dot{B}_{\infty}^{X\setminus\Lambda}$. For each $k\in\{1,\ldots,N_{X}\}$ we have
$[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{X}(\varphi^{k}_{\Lambda}\otimes\varphi^{k}_{X\setminus\Lambda})]$ $=[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{\Lambda}(\varphi^{k}_{\Lambda})]\otimes Q_{j}^{X\setminus\Lambda}(\varphi^{k}_{X\setminus\Lambda})$
$Q_{j}^{X}\big{(}\{a_{\Lambda},\varphi^{k}_{\Lambda}\otimes\varphi^{k}_{X\setminus\Lambda}\}_{B_{\infty}^{\Lambda}}\big{)}$ $=Q_{j}^{\Lambda}\big{(}\{a_{\Lambda},\varphi^{k}_{\Lambda}\}_{B_{\infty}^{\Lambda}}\big{)}\otimes Q_{j}^{X\setminus\Lambda}(\varphi^{k}_{X\setminus\Lambda})\,.$
If follows that
$\bigg{\|}\frac{2j+1}{i}[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{X}(\varphi^{k}_{\Lambda}\otimes\varphi^{k}_{X\setminus\Lambda})]-Q_{j}^{\Lambda}\big{(}\{a_{\Lambda},\varphi^{k}_{\Lambda}\otimes\varphi^{k}_{X\setminus\Lambda}\}_{B_{\infty}^{\Lambda}}\big{)}\bigg{\|}_{B_{j}^{\Lambda}}\\ \leq\bigg{\|}\frac{2j+1}{i}[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{\Lambda}(\varphi^{k}_{\Lambda})]-Q_{j}^{\Lambda}\big{(}\{a_{\Lambda},\varphi^{k}_{\Lambda}\}_{B_{\infty}^{\Lambda}}\big{)}\bigg{\|}_{B_{j}^{\Lambda}}\|\varphi^{k}_{X\setminus\Lambda}\|_{B_{\infty}^{X\setminus\Lambda}}\,.$
At this stage we consider the linear operator
$T_{j}\colon C^{1}(\mathbb{S}_{\Lambda})\to\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}\,,\\ [T_{j}(b_{\Lambda})]_{p}:=\begin{dcases}\frac{2j+1}{i}[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{\Lambda}(b_{\Lambda})]-Q_{j}^{\Lambda}\big{(}\{a_{\Lambda},b_{\Lambda}\}_{B_{\infty}^{\Lambda}}\big{)}&p=j\\ 0&p\neq j\end{dcases}\,,$
@@
where $\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}$ denotes the full $C^{*}$-direct product, cf. [10]. By direct inspection $T_{j}$ is linear and bounded, moreover, the DGR condition (3) implies that
@@
$\sup_{j\in\mathbb{Z}_{+}/2}\big{\|}T_{j}(b_{\Lambda})\big{\|}_{\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}}=\sup_{j\in\mathbb{Z}_{+}/2}\sup_{p\in\mathbb{Z}_{+}/2}\big{\|}[T_{j}(b_{\Lambda})]_{p}\big{\|}_{B_{p}^{\Lambda}}=\sup_{j\in\mathbb{Z}_{+}/2}\big{\|}[T_{j}(b_{\Lambda})]_{j}\big{\|}_{B_{j}^{\Lambda}}<\infty\,,$
for all $b_{\Lambda}\in C^{1}(\mathbb{S}_{\Lambda})$. By Banach-Steinhaus Theorem it follows that
$\sup_{j\in\mathbb{Z}_{+}/2}\big{\|}T_{j}\big{\|}_{C^{1}(\mathbb{S}^{2}_{\Lambda})\to\prod_{j\in\overline{\mathbb{Z}_{+}/2}}B_{j}^{\Lambda}}<\infty\,,$
therefore, for all $a_{\Lambda}\in\dot{B}_{\infty}^{\Lambda}$ there exists $C=C(a_{\Lambda})>0$ such that
$\bigg{\|}\frac{2j+1}{i}[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{\Lambda}\tilde{a}_{\Lambda}]-Q_{j}^{\Lambda}\big{(}\{a_{\Lambda},\tilde{a}_{\Lambda}\}_{B_{\infty}^{\Lambda}}\big{)}\bigg{\|}_{B_{j}^{\Lambda}}\leq C\|\tilde{a}_{\Lambda}\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})}\qquad\forall j\in\mathbb{Z}_{+}/2\,,\,\forall\tilde{a}_{\Lambda}\in\dot{B}_{\infty}^{\Lambda}\,.$
This implies
$\bigg{\|}\frac{2j+1}{i}[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{X}(\varphi_{X})]-Q_{j}^{X}\big{(}\{a_{\Lambda},\varphi_{X}\}_{B_{\infty}^{\Lambda}}\big{)}\bigg{\|}_{B_{j}^{X}}\leq C\sum_{k=1}^{N_{X}}|c_{k,X}|\|\varphi^{k}_{\Lambda}\|_{C^{1}(\mathbb{S}_{\Lambda})}{\|\varphi^{k}_{X\setminus\Lambda}\|_{B_{\infty}^{X\setminus\Lambda}}}\,.$
Since this holds for all decompositions of $\varphi_{X}$ we obtain
$\bigg{\|}\frac{2j+1}{i}[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{X}(\varphi_{X})]-Q_{j}^{X}\big{(}\{a_{\Lambda},\varphi_{\Lambda}\}_{B_{\infty}^{\Lambda}}\big{)}\bigg{\|}_{B_{j}^{X}}\leq C\|\varphi_{X}\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})\otimes B_{\infty}^{X\setminus\Lambda}}\leq C\|\varphi_{X}\|_{C^{1}(\mathbb{S}^{2}_{X})}\,.$
On account of condition (35) the series (44) converges. Moreover, a direct application of dominated converges shows that
$\lim_{j\to\infty}\sum_{X\cap\Lambda\neq\varnothing}\bigg{\|}\frac{2j+1}{i}[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{X}(\varphi_{X})]-Q_{j}^{\Lambda\cup X}\big{(}\{a_{\Lambda},\varphi_{X}\}_{B_{\infty}^{\Lambda}}\big{)}\bigg{\|}_{B_{j}^{\Lambda\cup X}}=0\,,$
proving (43). ∎

### v2

@@ 1 Introduction
In this framework $j\in\mathbb{Z}_{+}/2$ is interpreted as a semiclassical parameter —in fact, $h_{j}:=1/(2j+1)$ is the proper semiclassical parameter— and $j\to\infty$ corresponds to the semiclassical limit. Given a sequence $(\omega_{j})_{j\in\mathbb{Z}_{+}/2}$ of states such that $\omega_{j}\in S(\mathfrak{A}_{j})$, the semiclassical limit is obtained considering the weak*-limit points of the sequence $(\omega_{j}\circ Q_{j})_{j\in\mathbb{Z}_{+}/2}$ of functionals over $\mathfrak{A}_{\infty}$ —each such weak*-limit point defines a state on $\mathfrak{A}_{\infty}$.
@@ 2.2 Berezin SDQ for a single site system
identify a SDQ over a suitably defined bundle of $C^{*}$-algebras $\boxed{B_{\ast}}\subseteq\prod_{j\in\overline{\mathbb{Z}_{+}/2}}B_{j}$, cf. [37, Thm. 8.1]. From a physical point of view, the semiclassical parameter is identified with $\boxed{h_{j}}:=(2j+1)^{-1}$.
@@
where $P_{\ell,m}$ denotes the Legendre polynomial of order $\ell,m$. Notice that this choice of normalization is such that $\|Y_{\ell,m}\|_{L^{2}(\mathbb{S}^{2},\mu_{\ell})}=1$, moreover, $\|Y_{\ell,m}\|_{B_{\infty}}\leq 1$, cf. [51, Cor. 2.9]. The set $\{Y_{\ell,m}\}_{\ell,m}$ is a complete orthogonal system for $L^{2}(\mathbb{S}^{2},\mu_{0})$ made by orthogonal but not $L^{2}$-normalized vectors. With this convention we also have [42, §3.6.2]
@@ 2.3 Berezin SDQ on $\Gamma$
$\big{\|}Q_{j}^{\Gamma}\big{(}\{a_{\Lambda},\tilde{a}_{\Lambda}\}_{B_{\infty}^{\Gamma}}\big{)}-ih_{j}^{-1}\big{[}Q_{j}^{\Gamma}(a_{\Lambda}),Q_{j}^{\Gamma}(\tilde{a}_{\Lambda})\big{]}\big{\|}_{B_{j}^{\Gamma}}\\ =\big{\|}Q_{j}^{\Lambda}\big{(}\{a_{\Lambda},\tilde{a}_{\Lambda}\}_{B_{\infty}^{\Lambda}}\big{)}-ih_{j}^{-1}[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{\Lambda}\tilde{a}_{\Lambda}]\big{\|}_{B_{j}^{\Lambda}}\underset{j\to\infty}{\longrightarrow}0\,,$
@@ 3.2 Semi-classical limit of quantum KMS condition
The proof requires two main ingredients. The first one is a useful characterization of the $(\beta,\delta_{\infty}^{\Gamma})$-KMS classical condition, cf. Lemma 3.2: The latter is inspired by an analogous characterization of $(\beta,\delta_{j}^{\Gamma})$-KMS quantum states, cf. Remark 3.3. The second piece of information is a control of the semiclassical limit of the derivation $\delta_{j}^{\Gamma}$. Specifically, for any finite region $\Lambda\Subset\Gamma$, the DGR condition (3) implies that $h_{j}^{-1}\delta_{j}^{\Lambda}\circ Q_{j}^{\Lambda}-Q_{j}^{\Lambda}\circ\delta_{j}^{\Lambda}\underset{j\to\infty}{\longrightarrow}0$ strongly on $\dot{B}_{\infty}^{\Lambda}$. This is a key property which notably holds also for the lattice system on the entire $\Gamma$, cf. Lemma 3.4.
@@
$\lim_{j\to\infty}\Big{\|}h_{j}^{-1}\delta_{j}^{\Lambda}(Q_{j}^{\Lambda}(a_{\Lambda}))-Q_{j}^{\Lambda}(\delta_{\infty}^{\Lambda}(a_{\Lambda}))\Big{\|}_{B_{j}^{\Lambda}}=0\qquad\forall a_{\Lambda}\in\dot{B}_{\infty}^{\Lambda}\,.$ (42)
@@
$\lim_{j\to\infty}\Big{\|}h_{j}^{-1}\delta_{j}^{\Gamma}(Q_{j}^{\Gamma}(a_{\Lambda}))-Q_{j}^{\Gamma}(\delta_{\infty}^{\Gamma}(a_{\Lambda}))\Big{\|}_{B_{j}^{\Gamma}}=0\,.$ (43)
@@
$\Big{\|}h_{j}^{-1}\delta_{j}^{\Gamma}\big{(}Q_{j}^{\Gamma}(a_{\Lambda})\big{)}-Q_{j}^{\Gamma}\big{(}\delta_{\infty}^{\Gamma}(a_{\Lambda})\big{)}\Big{\|}_{B_{j}^{\Gamma}}\\ \leq\sum_{X\cap\Lambda\neq\varnothing}\Big{\|}\frac{1}{ih_{j}}[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{X}(\varphi_{X})]-Q_{j}^{\Lambda\cup X}\big{(}\{a_{\Lambda},\varphi_{X}\}_{B_{\infty}^{\Lambda}}\big{)}\Big{\|}_{B_{j}^{\Lambda\cup X}}\,.$
For each $j\in\mathbb{Z}_{+}/2$ the series
$\sum_{X\cap\Lambda\neq\varnothing}\Big{\|}\frac{1}{ih_{j}}[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{X}(\varphi_{X})]-Q_{j}^{\Lambda\cup X}\big{(}\{a_{\Lambda},\varphi_{X}\}_{B_{\infty}^{\Lambda}}\big{)}\Big{\|}_{B_{j}^{\Lambda\cup X}}\,,$
converges on account of condition (37). We will now prove that it vanishes in the limit $j\to\infty$. Notice that each term of the series vanishes as $j\to\infty$ on account of the DGR condition (3). We will now prove that each term can be bounded as
$\Big{\|}\frac{1}{ih_{j}}[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{X}(\varphi_{X})]-Q_{j}^{\Lambda\cup X}\big{(}\{a_{\Lambda},\varphi_{X}\}_{B_{\infty}^{\Lambda}}\big{)}\Big{\|}_{B_{j}^{\Lambda\cup X}}\leq c_{\Lambda}\|\varphi_{X}\|_{C^{1}(\mathbb{S}^{2}_{X})}\,,$
for a $j$-independent constant $c_{\Lambda}>0$. This will allows to apply dominated convergence, concluding the proof.
To this avail we consider the Banach space
$C^{0}(\mathbb{S}^{2}_{\Lambda^{c}},C^{1}(\mathbb{S}^{2}_{\Lambda}))=C^{1}(\mathbb{S}^{2}_{\Lambda})\otimes_{\varepsilon}B_{\infty}^{\Lambda^{c}}\,,\qquad\|f\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})\otimes_{\varepsilon}B_{\infty}^{\Lambda^{c}}}:=\sup_{\sigma_{\Lambda^{c}}\in\mathbb{S}^{2}_{\Lambda^{c}}}\|f(\sigma_{\Lambda^{c}})\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})}\,.$
where $\otimes_{\varepsilon}$ denotes the injective tensor product of Banach spaces, cf. [47, §3.2]. Let $T_{j}\colon\dot{B}_{\infty}^{\Gamma}\to\dot{B}_{j}^{\Gamma}$ be the linear operator defined by
$T_{j}(b_{X}):=\frac{1}{ih_{j}}[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{X}(b_{X})]-Q_{j}^{\Lambda\cup X}\big{(}\{a_{\Lambda},b_{X}\}_{B_{\infty}^{\Lambda}}\big{)}\,,$
for all $b_{X}\in\dot{B}_{\infty}^{X}\subset\dot{B}_{\infty}^{\Gamma}$, $X\Subset\Gamma$. By direct inspection we find
$\|T_{j}(b_{X})\|_{B_{j}^{\Gamma}}\leq c_{\Lambda}\|b_{X}\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})\otimes_{\varepsilon}B_{\infty}^{\Lambda^{c}}}\,,$
which implies that $T_{j}$ has a unique extension $T_{j}\colon C^{1}(\mathbb{S}^{2}_{\Lambda})\otimes_{\varepsilon}B_{\infty}^{\Lambda^{c}}\to B_{j}^{\Gamma}$. Notice that, the DGR condition (3) entails $\|T_{j}(b_{X})\|_{B_{j}^{\Gamma}}\underset{j\to\infty}{\longrightarrow}0$ for all $b_{X}\in\dot{B}_{\infty}^{X}$, $X\Subset\Gamma$. Moreover,
$\|T_{j}(b_{X})\|_{B_{j}^{\Gamma}}\leq\|T_{j}\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})\otimes_{\varepsilon}B_{\infty}^{\Lambda^{c}}\to B_{j}^{\Gamma}}\|b_{X}\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})\otimes_{\varepsilon}B_{\infty}^{\Lambda^{c}}}\leq\|T_{j}\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})\otimes_{\varepsilon}B_{\infty}^{\Lambda^{c}}\to B_{j}^{\Gamma}}\|b_{X}\|_{C^{1}(\mathbb{S}^{2}_{X})}\,,$
where $\|T_{j}\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})\otimes_{\varepsilon}B_{\infty}^{\Lambda^{c}}\to B_{j}^{\Gamma}}$ denotes the operator norm of $T_{j}$. Thus, if we were able to prove that $\sup_{j}\|T_{j}\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})\otimes_{\varepsilon}B_{\infty}^{\Lambda^{c}}\to B_{j}^{\Gamma}}<\infty$ then condition (37) would entail that
$\sum_{X\cap\Lambda\neq\varnothing}\Big{\|}\frac{1}{ih_{j}}[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{X}(\varphi_{X})]-Q_{j}^{\Lambda\cup X}\big{(}\{a_{\Lambda},\varphi_{X}\}_{B_{\infty}^{\Lambda}}\big{)}\Big{\|}_{B_{j}^{\Lambda\cup X}}=\sum_{X\cap\Lambda\neq\varnothing}\|T_{j}(\varphi_{X})\|_{B_{j}^{\Lambda}}\,,$
converges, moreover, it vanishes as $j\to\infty$ by dominated convergence, concluding the proof.
Thus, we prove that $\sup_{j}\|T_{j}\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})\otimes_{\varepsilon}B_{\infty}^{\Lambda^{c}}\to B_{j}^{\Gamma}}<\infty$. For that let us consider the linear map
$T_{j}^{\Lambda}\colon C^{1}(\mathbb{S}^{2}_{\Lambda})\to B_{j}^{\Lambda}\qquad T_{j}^{\Lambda}(b_{\Lambda}):=\frac{1}{ih_{j}}[Q_{j}^{\Lambda}(a_{\Lambda}),Q_{j}^{\Lambda}(b_{\Lambda})]-Q_{j}^{\Lambda}\big{(}\{a_{\Lambda},b_{\Lambda}\}_{B_{\infty}^{\Lambda}}\big{)}\,.$
We then “lift” $T_{j}^{\Lambda}$ to
$\hat{T}_{j}^{\Lambda}\colon C^{1}(\mathbb{S}^{2}_{\Lambda})\to\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}\,,\qquad[\hat{T}_{j}^{\Lambda}(b_{\Lambda})]_{p}:=\begin{dcases}T_{j}^{\Lambda}(b_{\Lambda})&p=j\\ 0&p\neq j\end{dcases}\,,$
@@
where $\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}$ denotes the full $C^{*}$-direct product, cf. [10]. By direct inspection $\hat{T}_{j}^{\Lambda}$ is linear and bounded, moreover, the DGR condition (3) implies that
$\sup_{j\in\mathbb{Z}_{+}/2}\big{\|}\hat{T}_{j}^{\Lambda}(b_{\Lambda})\big{\|}_{\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}}=\sup_{j\in\mathbb{Z}_{+}/2}\sup_{p\in\mathbb{Z}_{+}/2}\|[\hat{T}_{j}^{\Lambda}(b_{\Lambda})]_{p}\|_{B_{p}^{\Lambda}}=\sup_{j\in\mathbb{Z}_{+}/2}\|T_{j}^{\Lambda}(b_{\Lambda})\|_{B_{j}^{\Lambda}}<\infty\,,$
for all $b_{\Lambda}\in C^{1}(\mathbb{S}^{2}_{\Lambda})$. By Banach-Steinhaus Theorem it follows that
$\sup_{j\in\mathbb{Z}_{+}/2}\big{\|}\hat{T}_{j}^{\Lambda}\big{\|}_{C^{1}(\mathbb{S}^{2}_{\Lambda})\to\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}}<\infty\,.$
We then consider the inductive tensor product of $\hat{T}_{j}^{\Lambda}\colon C^{1}(\mathbb{S}^{2}_{\Lambda})\to\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}$ with $Q_{j}^{\Lambda^{c}}\colon B_{\infty}^{\Lambda^{c}}\to B_{j}^{\Lambda^{c}}$, cf. [47, Prop. 3.2]. The latter is the unique bounded linear map
$\hat{T}_{j}^{\Lambda}\otimes_{\varepsilon}Q_{j}^{\Lambda^{c}}\colon C^{1}(\mathbb{S}^{2}_{\Lambda})\otimes_{\varepsilon}B_{\infty}^{\Lambda^{c}}\to\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}\otimes B_{j}^{\Lambda^{c}}\,,$
which extends the algebraic tensor product $\hat{T}_{j}^{\Lambda}\otimes Q_{j}^{\Lambda^{c}}$, that is,
$(\hat{T}_{j}^{\Lambda}\otimes_{\varepsilon}Q_{j}^{\Lambda^{c}})(b_{\Lambda}\otimes b_{\Lambda^{c}})=\hat{T}_{j}^{\Lambda}(b_{\Lambda})\otimes Q_{j}^{\Lambda^{c}}(b_{\Lambda^{c}})\,,$
for all $b_{\Lambda}\in C^{1}(\mathbb{S}^{2}_{\Lambda})$ and $b_{\Lambda^{c}}\in B_{\infty}^{\Lambda^{c}}$. Notice that $\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}\otimes B_{j}^{\Lambda^{c}}$ is a $C^{*}$-algebra once completed with its unique $C^{*}$-cross norm. The uniqueness of such norm is due to the fact that $B_{j}^{\Lambda^{c}}$ is the $C^{*}$-inductive limit of finite dimensional, hence nuclear, $C^{*}$-algebras, therefore, it is nuclear as well, cf. [10, §II.9.4.5]. We recall that
$\|\hat{T}_{j}^{\Lambda}\otimes_{\varepsilon}Q_{j}^{\Lambda^{c}}\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})\otimes_{\varepsilon}B_{\infty}^{\Lambda^{c}}\to\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}\otimes B_{j}^{\Lambda^{c}}}$ $=\|\hat{T}_{j}^{\Lambda}\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})\to\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}}\|Q_{j}^{\Lambda^{c}}\|_{B_{\infty}^{\Lambda^{c}}\to B_{j}^{\Lambda^{c}}}$
$=\|\hat{T}_{j}^{\Lambda}\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})\to\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}}\,,$
@@
therefore, $\sup_{j}\|\hat{T}_{j}^{\Lambda}\otimes_{\varepsilon}Q_{j}^{\Lambda^{c}}\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})\otimes_{\varepsilon}B_{\infty}^{\Lambda^{c}}\to\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}\otimes B_{j}^{\Lambda^{c}}}<\infty$. Finally, let
$\hat{T}_{j}\colon C^{1}(\mathbb{S}^{2}_{\Lambda})\otimes_{\varepsilon}B_{\infty}^{\Lambda^{c}}\to\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}\otimes B_{j}^{\Lambda^{c}}\qquad[\hat{T}_{j}(f)]_{p}:=\begin{cases}T_{j}(f)&p=j\\ 0&p\neq j\end{cases}\,.$
Then $\hat{T}_{j}$ is linear and continuous, moreover, by direct inspection
$\hat{T}_{j}(b_{\Lambda}\otimes b_{\Lambda^{c}})=\hat{T}_{j}^{\Lambda}(b_{\Lambda})\otimes Q_{j}^{\Lambda^{c}}(b_{\Lambda^{c}})\,.$
It follows that $\hat{T}_{j}=\hat{T}_{j}^{\Lambda}\otimes_{\varepsilon}Q_{j}^{\Lambda^{c}}$ and therefore
$\|T_{j}\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})\otimes_{\varepsilon}B_{\infty}^{\Lambda^{c}}\to B_{j}^{\Gamma}}$ $=\|\hat{T}_{j}\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})\otimes_{\varepsilon}B_{\infty}^{\Lambda^{c}}\to\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}\otimes B_{j}^{\Lambda^{c}}}$
$=\|\hat{T}_{j}^{\Lambda}\otimes_{\varepsilon}Q_{j}^{\Lambda^{c}}\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})\otimes_{\varepsilon}B_{\infty}^{\Lambda^{c}}\to\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}\otimes B_{j}^{\Lambda^{c}}}$
$=\|\hat{T}_{j}^{\Lambda}\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})\to\prod_{p\in\mathbb{Z}_{+}/2}B_{p}^{\Lambda}}\,,$
showing that $\sup_{j}\|T_{j}\|_{C^{1}(\mathbb{S}^{2}_{\Lambda})\otimes_{\varepsilon}B_{\infty}^{\Lambda^{c}}}<\infty$.